The half-life of a first order reaction is 100 seconds. What is the time required for 90% completion of the reaction?

Grade:8

1 Answers

Nikhil
6 years ago
Rate constant for this first ordered reaction is K=0.693/100Now , Time required for 90%completion is T=2.303/0.00693 ×log100/100-90 T=2.303/0.00693×log 10 T=332.2259 second